OpenStudy (anonymous):

If x+y=15 and x-y=25, then what is the value of 2x^2 - y^2? A) 425 B) 775 C) 825 D) 1,575 E) 1,625

OpenStudy (anonymous):

a couple of ways to go about this. Less interesting is the straight forward way of solving for x and y by adding the equations together to get: \[2x = 40 \iff x = 20\] Then y will be: \[20+y = 15 \iff y = -5\] plugg and chug after that.

OpenStudy (anonymous):

\[(x-y)(x+y) = x^2-y^2\]

OpenStudy (anonymous):

@alchemista it doesnt work that y :o

OpenStudy (anonymous):

it does <.< his way is more interesting. you multiply, then add on an x^2.

OpenStudy (angela210793):

@Alchemista is just x^2 multiplied with 2 not all tht

OpenStudy (anonymous):

i did that way but got stuck when i see the 2x^2

OpenStudy (anonymous):

\[2x^2-y^2 = x^2+x^2-y^2 = x^2+(x+y)(x-y)\]

OpenStudy (anonymous):

I misread it, yes indeed its not all multiplied by 2

OpenStudy (anonymous):

ohh i see what you guys try to say ^.^ thank you all xD